A torque of 2 N.m increases a wheel's angular speed by 3 times in 4.0 s. The rotational inertia of the wheel is 0.04 kg.m^2. The initial angular speed of the wheel is
1
Expert's answer
2019-06-12T11:23:11-0400
"\tau=I\alpha"

"\alpha=\frac{3\omega-\omega}{t}=\frac{2\omega}{t}"

"\omega=\frac{\tau t}{2I}=\frac{(2)(4)}{2(0.04)}=100\frac{rad}{s}"